OpenStudy (anonymous):

Evaluate 6C4 2 30 24 15

OpenStudy (saifoo.khan):

15

OpenStudy (anonymous):

\[\frac{6\times 5}{2}=15\]

OpenStudy (anonymous):

so you too 6*5 then divided by 2 gets your answer?

OpenStudy (amistre64):

\[\frac{6.5.4.3.2.1}{(4.3.2.1)(2.1)}=\frac{6!}{(6-2)!\ 2!}=15\]